how do you solve this.. x+2y=5 solve each system of equation
x-y=-1

Answer by ReadingBoosters(3246)

x + 2y = 5
x - y = -1
Subtract the equations
x - x + 2y - -y = 5 - -1
3y = 6
3y/3 = 6/3
y = 2
Multiply second eqution by 2
2(x-y=-1)
2x-2y=-2
Add to the first equation
x +2x +2y + -2y = 5 + -2
3x = 3
3x/3=3/3
x=1
Therefore,
x=1, y = 2
Proof
1 + 2(2) = 5
1-2=-1